OpenStudy (anonymous):

Can someone help me with this? A magician is doing a rope trick with a rope that is 7/8ft long. During the trick, he cuts a piece that is 1/6 of the original. How long is the cut piece?

OpenStudy (anonymous):

\[\frac{7}{8}\times \frac{1}{6}\]

OpenStudy (anonymous):

I still dont understand.